OpenStudy (insa):

solve by first converting to quadratic equation 8/x- 3/x+1=5

OpenStudy (abmon98):

subtract 8/x-3/x+1=5 5/x+1=5 get rid of fraction by multiplying by x 5+x=5x 4x=5 x=5/4